OpenStudy (anonymous):

How do I solve this? Simplify square root of 72 end root minus square root of 147 end root minus 5 square root of 8 end root plus 3 square root of 48. Thank you so much!

OpenStudy (anonymous):

sqrt(72) - sqrt(147) - 5sqrt(8) + 3sqrt(48) = sqrt(9 * 4 * 2) - sqrt(49 * 3) - 5sqrt(4 * 2) + 3sqrt(16 * 3) = 3*2*sqrt(2) - 7sqrt(3) - 2*5sqrt(2) + 4*3sqrt(3) = 6sqrt(2) - 10sqrt(2) - 7sqrt(3) + 12sqrt(3) = -4sqrt(2) + 5sqrt(3)

OpenStudy (anonymous):

@chitters Thank you! :)